### v2

# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

Abstract
Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes. Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders. In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high performance and rapid inference speeds in the domain of STR. Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder. This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes. Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR. Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds. Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed. Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition. The code is publicly available at https://github.com/cxfyxl/VIPTR.
1 Introduction
Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al. (2016); Guo et al. (2022); Yang et al. (2020); Guo et al. (2023); Zhou et al. (2023); Li et al. (2023a); Wu et al. (2023). Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks. Despite progress, STR still faces hurdles such as distorted text, varied fonts, shadows, and complex backgrounds. Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
Figure 1: Model Architecture Evolution for Scene Text Recognition.
Scene text recognition initially relied on deep learning models like CRNNs Shi et al. (2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances. Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds. More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise. Recent researchers 2021 (2020); Zhang et al. (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss. From the perspective of simplifying the model architecture, a number of models used only CNN Borisyuk et al. (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods. Inspired by the pyramid structure of Swin Transformer Liu et al. (2021), a SVTR model Du et al. (2022) with single visual modal input only is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information. SVTR has four structural variants with different capacities.
The STR models based on vision Transformer set a typical example for fast and effective STR schemes for single visual modality. On one hand, it practices the effectiveness of replacing convolution with self-attention mechanism to realize a new visual feature extractor. On the other hand, it reveal the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs. Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model. Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al. (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background. This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic relationships between different characters. In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al. (2021) combined with manhattan self-attention (MaSA) Fan et al. (2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al. (2023) module to model global dependencies. Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing. Compared with SWin, this method significantly reduces the time complexity and does not require the use of a fixed-size mask group to limit the length of the input image. At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images. To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
The main contributions of this work can be summarized as.
  • We verified that the single-vision model based on the self-attention mechanism can still achieve comparable accuracy to the high-level vision-language model in the scene text recognition task. At the same time, the use of sparse operators and different attention combinations can accelerate the calculation of the self-attention mechanism, and achieve the balance between performance and speed advantages that STR has been pursuing.
  • We proposed VIPTR, a feature extraction module tailored for parsing text. The module can adapt to and accurately recognize the text image input of any length and has the versatility of cross-language recognition, which has a good application prospect.
  • We verified the superiority of VIPTR on cross-language benchmark datasets and manually annotated industrial application datasets. Among them, VIPTRv1-L achieved SOTA in recognizing Chinese and English scene text. VIPTRv2-T achieves the most efficient inference while ensuring accuracy. It is inferred on an NVIDIA V100 GPU with a parameter of 5.1M and consumes an average of 3.3ms per text image.
2 Related Work
Figure 2: Overall architecture of the proposed VIPTR. It is a four-stage network with three-stage height progressively decreased. In each stage, a series of attention mixing blocks are carried out and followed by a subsampling or combining operation. In particular, VIPTR designs two visual feature fusion modes in vision permulation blocks: series and parallel. At last, the recognition is conducted by CTC decoder.
CNN-Based Vision Modules
Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks. VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model. Later, ResNet He et al. (2016) introduced residual connections along with the batch normalization (BN) layer, which enabled training of exceedingly deep neural networks and further improved performance. In the realm of STR, the CNN-based vision model undergoes a transformation into a pyramid structure capable of processing local features across multiple stages. This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN). High recognition accuracy can also be obtained, which proves the effectiveness of locality bias extracted by CNN. We believe that the experience gained in CNN-based vision models, including the pyramid structure for extracting locality bias through small kernel convolutions, residual connections, and step-by-step downsampling, remains valid in current research. Therefore, these techniques continue to be utilized in our proposed STR architecture.
Transformer-Based Vision Modules
ViT Dosovitskiy et al. (2020) is the first work to introduce Transformer Vaswani et al. (2017); Yang et al. (2022); Touvron et al. (2021) into the construction backbone of vision tasks. Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks. DeiT Touvron et al. (2021) introduces a new training strategy, enabling ViT to achieve superior performance compared to CNNs on small-scale training sets such as ImageNet-1K. These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms. Later, the pyramid structure also enhances the performance of Transformer-based models across various vision tasks. Based on the pyramid structure, Swin Transformer proposes to use self-attention within each local window. This can be regarded as a use of locality bias. Currently, vision Transformers are integrated into STR task as visual feature encoders, leading to a significant improvement in accuracy. ViTSTR Atienza (2021), visionLAN Wang et al. (2021b), and MaskOCR Lyu et al. (2022) have all come close to the SOTA effect. Among them, SVTR-T is in the context of inputting constant-width text images, it both ensures high performance and creates the fastest recognition efficiency. Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens. As the resolution of the input image increases, the inference time consumption also increases quadratically. square growth.
Efficient Self-Attention Based Vision Modules
A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al. (2017); Yang et al. (2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining. The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs. Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale. Spatial reduction attention (SRA) Wang et al. (2021a) effectively extracts global information by utilizing sparse token-region relationships, and OSRA introduces overlapping spatial reduction (OSR) Lou et al. (2023) into SRA to achieve better performance by using larger and overlapping patches.
3 Method
3.1 Design Guidelines
In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high performance networks for STR. Here are the design guidelines we followed:
  1. 1.
    Explicitly inject character features into the front of the network in the form of locality bias.
  2. 2.
    Design a new permutation method to effectively fuse the character component features of the Local and Global Attention modules in the middle layer.
  3. 3.
    Implicitly handle image aspect and width resolution in the network initialization structure, canceling or replacing all modules that depend on the input size.
3.2 Overall Architecture
The overview of the proposed VIPTR is illustrated in Figure 2. It is a four-stage network with three height progressively decreased layers dedicated to text recognition. For an image text of size $H\times W\times 3$, it is first transformed to $\frac{H}{4}\times\frac{W}{4}$ patches of channel dimension $C_{0}$ via a two-layer CNN-based patch embedding. Since each patch is associated with a text character or part of the background in the image, these patches are called character components. Then, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension. Local and Global Attention blocks are used to focus on character feature extraction and dependency capture between character components. The backbone is used to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which perceives multi-granularity character features. Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
3.3 Patch Embedding
The first step of vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences. In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$. Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs. This schemeDu et al. (2022) gradually increases feature dimensions, which is beneficial to feature fusion.
Figure 3: Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
3.4 Character and String Modeling
End-to-end text recognizers rely heavily on the mining of character features and string context. However, most of the existing studies use feature sequences with height $1$ to represent image text. Each feature corresponds to a thin image region, which is usually noisy, especially for irregular text. This is not the best way to describe the characters of an image. The vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds. This shows that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual relationships of strings.
More specifically, image text requires features to be extracted from two perspectives. The first is local character features, such as strokes and fonts. It encodes morphological features and correlations between different parts of a character. The second is inter-character context relations, such as correlations between different characters or between text and non-text components. Therefore, we design two classes of mixing blocks to sense relevance by using self-attention with different receptive fields. Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features. When attention mixing block is input, the Layer Normalization (LN), Position Embedding (PE) are applied and then the self-attention mechanism is performed for dependency modeling. Next, LN and Feed Forward Network (FFN) are successively applied for feature fusion. Finally, the residual connection is used.
Local Attention.
The goal of Local Attention is to encode character morphological features and establish correlations within character components, thereby simulating stroke features. Local Attention needs to consider the neighborhood of character components on the two-dimensional coordinate system. Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system. MaSA is shown in Figure 3LABEL:sub@subfig:dmsa. The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened. CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin. This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel. For image text input, considering that the image resolution is larger than Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing and the length and width resolution needs to be guaranteed. Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
Global Attention.
As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual relationships between all character components. Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters. In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components. In addition, in order to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information. The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
Positional Embedding.
Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back. Typical positional Embedding mechanisms include absolute positional encoding (APE)Vaswani et al. (2017), relative positional encoding (RPE) Liu et al. (2021) and CPE. APE and RPE are designed for a specific input size and are not friendly to varying input resolutions. CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions. Then the generated positional encoding will be added onto the input feature. Locally-enhanced Positional Encoding (LePE)Dong et al. (2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
As mentioned above, the two types of hybrid blocks aim to extract complementary different features. In VIPTR, blocks are processed locally first and then globally. At each stage, Local and Global Attentions are recurrently applied multiple times for comprehensive feature extraction. Effect of the two kinds of blocks arrangements will be shown in next ablation experiments.
3.5 Subsampling
In order to save computational costs and extract significant features layer by layer, we added a down-sampling operationDu et al. (2022) on the height dimension after the first and third attention mixing stages of VIPTR. Using the features output from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. The Subsampling block halves the height while maintaining a constant width. It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the information loss.
3.6 Combining and Prediction
In the final stage, we use combinatorial operationsDu et al. (2022) to integrate sequence features. It first uses AdaptiveAvgPool to compress the height dimension to $1$, followed by fully connected layers, Hardswish activation, and dropout. By doing so, the character components are further compressed into feature sequences, where each element is represented by a feature of length $C_{3}$. Using AdaptiveAvgPool avoids applying convolution to very small embeddings of one dimension, such as height $2$, while being able to apply a variety of initial input heights.
We finally use the CTC decoder to perform parallel linear prediction on the obtained combined features to achieve recognition. Specifically, a linear classifier with $N$ nodes is adopted. It generates transcribed sequences of size $W/4$, where components of identical characters are ideally transcribed as repeating characters and components of non-text are transcribed as whitespace symbols. The sequence is automatically compressed into the final result. For example, $N$ is set to $37$ for English when implemented.
3.7 Architecture Variants
Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2. Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, number of mixing blocks. By varying them, VIPTR architectures with different number of parameters can be obtained. In summary, we constructed four typical architectures, namely VIPTRv1-T(Tiny), VIPTRv2-T, VIPTRv1-L(Large) and VIPTRv2-B(Base). Its detailed configuration is shown in Table 2.
Atten. Permutation IC13 IC15 FLOPs (G)
CRNN(None) 75.3 56.7 0.69
SVTR-T 77.9 62.5 0.29
$[L_{1}][L_{1}G_{1}][G_{1}]$ 79.1 65 0.269
$[L_{1}][L_{1}G_{2}][G_{1}]$ 79.1 65.2 0.258
$[L_{2}][L_{2}G_{1}][G_{1}]$ 79 65.1 0.282
$[L_{2}][L_{2}G_{2}][G_{1}]$ 79.3 65.3 0.27
$[L_{1}][L_{2}G_{1}][G_{1}]$ 78.8 64.9 0.27
$[L_{2}][L_{1}G_{1}][G_{1}]$ 78.9 65.1 0.27
$[L_{1}][L_{2}G_{2}][G_{1}]$ 78.7 65 0.258
$[L_{2}][L_{1}G_{2}][G_{1}]$ 78.8 64.8 0.258
$[L_{1}][L_{1}//G_{1}][G_{1}]$ 75.5 60.5 0.195
$[L_{1}][L_{1}//G_{2}][G_{1}]$ 75.4 60.5 0.192
$[L_{2}][L_{2}//G_{1}][G_{1}]$ 75.6 60.4 0.197
$[L_{2}][L_{2}//G_{2}][G_{1}]$ 75.8 61 0.194
$[L_{1}][L_{2}//G_{1}][G_{1}]$ 75 59.9 0.196
$[L_{2}][L_{1}//G_{1}][G_{1}]$ 75.2 59.7 0.196
$[L_{1}][L_{2}//G_{2}][G_{1}]$ 74.9 60.1 0.193
$[L_{2}][L_{1}//G_{2}][G_{1}]$ 75.2 60.3 0.193
Table 1: Ablation study on Attention Blocks’ permutations.
Model $[C_{0},C_{1},C_{2},C_{3}]$ $[N_{1},N_{2},N_{3},N_{4}]$ Heads Attention Permutation Params (M) FLOPs (G)
VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 3.99 0.26
VIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.26 0.19
VIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
Table 2: VIPTR architecture variants (w/o counting the rectification module and linear classifier).
4 Experiments
4.1 Datasets
For English STR task, our models are trained on two commonly used synthetic scene text datasets and a real scene dataset, i.e., MJSynth (MJ)Jaderberg et al. (2014, 2016) and SynthText (ST)Gupta et al. (2016) trainset (72k+55k), COCO-Text Gomez et al. (2017)(38062) and the trainsets union of ICDAR 2013 (IC13)Karatzas et al. (2013), ICDAR 2015 (IC15)Karatzas et al. (2015), IIIT5K Mishra et al. (2012) and Street View Text (SVT) Wang et al. (2011). Then the models are tested on six public benchmarks as follows: IC13 testset contains 857 selected regular testing images. SVT testset has 647 testing images cropped form Google Street View. IIIT5K-3000 is collected from the website and contains 3000 test images. IC15 is taken with Google Glasses without careful position and focusing, and testset we used has 1811 images. Street View Text Perspective (SVTP)Phan et al. (2013) is also cropped from Google Street View. There are 639 test images in this set. CUTE80 (CUTE) is proposed in Risnumawan et al. (2014) for curved text recognition. 288 testing images are cropped from full images by using annotated words.
For Chinese recognition task, we use the Chinese Scene and Document Dataset Yu et al. (2021). The former is a public dataset containing 509,164, 63,645 and 63,646 training, validation, and test images. A total of 5,880 different characters appear. The latter is also a public dataset, which containing 400,000, 50,000 and 50,000 training, validation, and test images. A total of 4,865 different characters appear. The validation set is utilized to determine the best model, which is then assessed using the test set.
Positional Embedding IC13 IC15 FLOPs(G)
None 76.7 59.7 0.25
APE 77.3 63.5 0.26
CPE 77.3 63.6 0.26
LePE 79.1 65 0.26
Table 3: Ablation study on Positional Embeddings.
Model Testset1 Testset2 Testset3
CRNN 97.5 91 91.6
SVTR-T 98.2 47.5 5.9
VIPTRv1-T 98.8 92.1 93.7
Table 4: Ablation studies on inputs of different lengths.
4.2 Implementation Details
VIPTR uses the rectification module Shi et al. (2018), where the image text is resized to $32\times 64$ for distortion correction. We use the AdamW optimizer with weight decay of $0.05$ for training. For English models, when CSWin is used by Local Attention, the input image size is set to $32\times 96$. Otherwise, it is set to $32\times 100$. The initial learning rate are set to $\frac{5}{10^{4}}\times\frac{batchsize}{2048}$. The cosine learning rate scheduler with $2$ epochs linear warm-up is used in all $49$ epochs. Data augmentation like stretch, perspective distortion, blur and Gaussian noise, are randomly performed during the training with probability $0.4$. The alphabet includes all case-insensitive alphanumerics. The maximum annotation length corresponding to the training sample is set to $25$. For Chinese models, the input size is set to $32\times 320$ and all images are resized or padded to this scale during training. The initial learning rate are set to $\frac{3}{10^{4}}\times\frac{batchsize}{512}$. The cosine learning rate scheduler with $5$ epochs linear warm-up is used in all $100$ epochs. Data augmentation was not used for training. Word accuracy is used as the evaluation metric. All models are trained by using 4 Tesla V100 GPUs on PyTorch.
4.3 Ablation Study
To better understand VIPTR, we perform controlled experiments on both IC13 (regular text) and IC15 (irregular text) under different settings. The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images and 500epochs of training. In addition, we also add aset of comparative experiments to verify the effectiveness of models on arbitrary lengths’ image texts. The Chinese Document Dataset is used for training and validation. Three test sets are prepared. Testset1 is the standard test set of Chinese Document Dataset, Testset2 (2367) is croped from real documents with image length in the range $[320,1500]$. Testset3 (638) gets from the same origin, whose image length is above 1500. For efficiency, all the experiments are carried out by using VIPTR-T without rectification module and data augmentation.
Methods English regular English unregular Avg Chinese Params Speed $\downarrow$
IC13 SVT IIIT5k IC15 SVTP CUTE Scene (M) (ms)
CRNNShi et al. (2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
RosettaBorisyuk et al. (2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
ViTSTRAtienza (2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
SVTR-TDu et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.03 4.5
SVTR-LDu et al. (2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
ASTERShi et al. (2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
NRTRSheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84 - 31.7 160
SARLi et al. (2019) 91 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120
SATRNLee et al. (2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
SRNYu et al. (2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
ABINetFang et al. (2021) 97.4 93.5 96.2 86 89.3 89.2 92.6 - 36.7 51.3
VisionLANWang et al. (2021b) 95.7 91.7 95.8 83.7 86 88.5 91.2 - 32.8 28
VSTTang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64 -
MATRNNa et al. (2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
MaskOCR*Lyu et al. (2022) 97.7 93.7 95.4 86.6 89 87.5 92.5 68.6 36 24.1
MaskOCRLyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100 -
TrOCRLi et al. (2023b) 98.3 93.2 91 84 91 89.6 90.2 - 558 20.8
AutoSTRZhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84 89.5 - 10.4 207
PREN*Yan et al. (2021) 94.7 92 92.1 79.2 83.9 81.3 88 - 29.1 40
VIPTRv1-T(Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.89 4.2
VIPTRv2-T 95.8 92.6 92.6 85.4 86 85 90.3 65.8 5.1 3.3
VIPTRv1-L(Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
VIPTRv2-B(Base) 97 94.2 94.8 88 90 90.2 92.7 71 22.1 8.0
Table 5: Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and others. * means lightweight version of the corresponding model. The speed is the average inference time on one NVIDIA Tesla V100 GPU with batchsize 4 (to align with Du et al. (2022)) for 3000 English image texts.
4.3.1 The Effective Permutation of Attention Blocks
There are various ways to match the global and Local Attention blocks in each stage. In previous researchvDu et al. (2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model. The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations. $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly. As can be seen, almost every scheme gives certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by self-attention mechanism. The relatively large gains on irregular text further explains the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
4.3.2 The Effectiveness of Positional Embeddings.
As seen in Table 3, different PEs behave slightly different in recognition accuracy. LePE scheme outperforms the default one (None) by 2.4% and 5.3% on the two datasets, indicating that it is effective especially for irregular text.
4.3.3 The Results for different lengths inputs testing
As seen in Table 4, the performance of the three models on the three testsets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. VIPTR not only supports arbitrary length of image text input, but the inference performance will not decrease as the input length increases.
4.4 Comparison with State-of-the-Art
We compare VIPTR with previous works on six English benchmarks covering regular and irregular text and one Chinese scene dataset in Table 5. We first look into the result in English datasets. Even VIPTRv2-T, the smallest one of the VIPTR family, achieves highly competitive accuracy among the w/o LM ones. When further compared with LM-used ones and SVTR-L, VIPTRv1-L gains the best accuracy on IC15 among the six English benchmarks. VIPTRv1-L and VIPTRv2-B’s overall accuracy are on par with recent studiesFang et al. (2021); Tang et al. (2021); Du et al. (2022); Lyu et al. (2022). As a contrast, VIPTRv1-L and VIPTRv2-B enjoy its simplicity and runs faster. We then analyze the result on Chinese Scene Dataset. Du et al. (2022) gives the accuracy of 8 existing methods as shown in the table. Encouragingly, VIPTR performs considerably well. Compared with SVTR-L, the best one among the listed ones, accuracy gains 0.5%. Other VIPTR methods also have advanced Chinese recognition accuracy. The above results are explained as VIPTRs can better perceive multi-grained character component features. They also achieved breakthroughs in characterizing Chinese words with rich stroke patterns.
5 Conclusion
We have presented VIPTR, a customized vision permutable extractor for fast and efficient STR. It extracts multi-grained character features that describe both stroke-like local patterns and inter-component dependence of varying distance at multiple height scales. Therefore, enjoying the merits of accuracy, efficiency and cross-lingual versatility. VIPTR with different capacities are also devised to meet diverse application needs. Experiments on both English and Chinese benchmarks basically verify the proposed VIPTR. Highly competitive or even better accuracy is observed compared to state-of-the-art methods, while running faster. We hope to combine VIPTR with an efficient text decoder to further improve the performance of STR.

### v3

@@
 Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders.
-In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high performance and rapid inference speeds in the domain of STR.
+In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high-performance and rapid inference speeds in the domain of STR.
 Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder.
@@
 (2020); Guo et al.
-(2023); Zhou et al.
 (2023); Li et al.
 (2023a); Wu et al.
+(2023); Zhou et al.
 (2023).
@@
 (2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss.
-From the perspective of simplifying the model architecture, a number of models used only CNN Borisyuk et al.
+From the perspective of simplifying the model architecture, some models used only CNN Borisyuk et al.
 (2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods.
 Inspired by the pyramid structure of Swin Transformer Liu et al.
-(2021), a SVTR model Du et al.
-(2022) with single visual modal input only is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information.
+(2021), the SVTR model Du et al.
+(2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information.
 SVTR has four structural variants with different capacities.
-The STR models based on vision Transformer set a typical example for fast and effective STR schemes for single visual modality.
-On one hand, it practices the effectiveness of replacing convolution with self-attention mechanism to realize a new visual feature extractor.
-On the other hand, it reveal the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs.
+The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality.
+On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor.
+On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs.
 Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
@@
 (2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background.
-This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic relationships between different characters.
+This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters.
 In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al.
@@
 Figure 2: Overall architecture of the proposed VIPTR.
-It is a four-stage network with three-stage height progressively decreased.
-In each stage, a series of attention mixing blocks are carried out and followed by a subsampling or combining operation.
-In particular, VIPTR designs two visual feature fusion modes in vision permulation blocks: series and parallel.
-At last, the recognition is conducted by CTC decoder.
+It is a four-stage network with three-stage height progressively decreasing.
+In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation.
+In particular, VIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel.
+At last, the recognition is conducted by the CTC decoder.
 CNN-Based Vision Modules
@@
 Later, ResNet He et al.
-(2016) introduced residual connections along with the batch normalization (BN) layer, which enabled training of exceedingly deep neural networks and further improved performance.
-In the realm of STR, the CNN-based vision model undergoes a transformation into a pyramid structure capable of processing local features across multiple stages.
+(2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance.
+In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages.
 This transformation effectively compresses the two-dimensional image input into a one-dimensional feature sequence, allowing for decoding the sequence without reliance on recurrent neural networks (RNN).
@@
 DeiT Touvron et al.
-(2021) introduces a new training strategy, enabling ViT to achieve superior performance compared to CNNs on small-scale training sets such as ImageNet-1K.
+(2021) introduces a new training strategy, enabling ViT to achieve superior performance on small-scale training sets such as ImageNet-1K.
 These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms.
@@
 This can be regarded as a use of locality bias.
-Currently, vision Transformers are integrated into STR task as visual feature encoders, leading to a significant improvement in accuracy.
+Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy.
 ViTSTR Atienza (2021), visionLAN Wang et al.
@@
 (2022) have all come close to the SOTA effect.
-Among them, SVTR-T is in the context of inputting constant-width text images, it both ensures high performance and creates the fastest recognition efficiency.
+SVTR-T both ensures high performance and creates the fastest recognition efficiency.
 Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens.
 As the resolution of the input image increases, the inference time consumption also increases quadratically.
-square growth.
 Efficient Self-Attention Based Vision Modules
@@
 Spatial reduction attention (SRA) Wang et al.
-(2021a) effectively extracts global information by utilizing sparse token-region relationships, and OSRA introduces overlapping spatial reduction (OSR) Lou et al.
+(2021a) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) Lou et al.
 (2023) into SRA to achieve better performance by using larger and overlapping patches.
@@
 3.1 Design Guidelines
-In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high performance networks for STR.
+In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high-performance networks for STR.
 Here are the design guidelines we followed:
@@
 3.2 Overall Architecture
-The overview of the proposed VIPTR is illustrated in Figure 2.
-It is a four-stage network with three height progressively decreased layers dedicated to text recognition.
-For an image text of size $H\times W\times 3$, it is first transformed to $\frac{H}{4}\times\frac{W}{4}$ patches of channel dimension $C_{0}$ via a two-layer CNN-based patch embedding.
+The schematic representation of the proposed VIPTR is depicted in Figure 2.
+It is a four-stage network with three height progressively decreased layers.
+Given an image text of size $H\times W\times 3$, it undergoes an initial transformation into $\frac{H}{4}\times\frac{W}{4}\times C_{0}$ patches through a two-layer CNN-based patch embedding.
 Since each patch is associated with a text character or part of the background in the image, these patches are called character components.
-Then, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension.
-Local and Global Attention blocks are used to focus on character feature extraction and dependency capture between character components.
-The backbone is used to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which perceives multi-granularity character features.
+Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension.
+Both Local and Global Attention blocks are employed to concentrate on character feature extraction and capture dependencies between character components.
+The extractor serves to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which captures multi-granularity character features.
 Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
 3.3 Patch Embedding
-The first step of vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences.
+The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences.
 In this article, for an image text, first convert it from $X\in R^{H\times W\times 3}$ into several character components $B_{0}B_{1}\in R^{\frac{H}{4}\times\frac{W}{4}\times C_{0}}$.
@@
 3.4 Character and String Modeling
-End-to-end text recognizers rely heavily on the mining of character features and string context.
-However, most of the existing studies use feature sequences with height $1$ to represent image text.
-Each feature corresponds to a thin image region, which is usually noisy, especially for irregular text.
-This is not the best way to describe the characters of an image.
-The vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds.
-This shows that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual relationships of strings.
+Existing studies mainly use feature sequences with height $1$ to represent image text.
+Each feature corresponds to a small local region.
+However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well.
+In contrast, the Vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds.
+This suggests that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual dependencies of strings.
 More specifically, image text requires features to be extracted from two perspectives.
-The first is local character features, such as strokes and fonts.
-It encodes morphological features and correlations between different parts of a character.
-The second is inter-character context relations, such as correlations between different characters or between text and non-text components.
-Therefore, we design two classes of mixing blocks to sense relevance by using self-attention with different receptive fields.
+The first involves local character features, including elements like strokes and fonts, representing internal associations of a character and morphological features.
+The second is context relations, such as correlations between different characters or between text and non-text components.
+Therefore, we design two classes of attention-mixing blocks with different receptive fields to sense relevance.
 Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features.
-When attention mixing block is input, the Layer Normalization (LN), Position Embedding (PE) are applied and then the self-attention mechanism is performed for dependency modeling.
-Next, LN and Feed Forward Network (FFN) are successively applied for feature fusion.
-Finally, the residual connection is used.
+Upon the input of the attention mixing block, Layer Normalization (LN) and Position Embedding (PE) are first applied.
+Subsequently, the self-attention mechanism is employed for dependency modeling.
+Finally, LN and Feed Forward Network (FFN) are successively applied for feature fusion, with the residual connection applied.
+Positional Embedding.
+Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back.
+Typical positional Embedding mechanisms include absolute positional encoding (APE) Vaswani et al.
+(2017), relative positional encoding (RPE) Liu et al.
+(2021), and CPE.
+APE and RPE are designed for a specific input size and are not friendly to varying input resolutions.
+CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions.
+Then the generated positional encoding will be added to the input feature.
+Locally-enhanced Positional Encoding (LePE) Dong et al.
+(2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
 Local Attention.
@@
 This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel.
-For image text input, considering that the image resolution is larger than Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing and the length and width resolution needs to be guaranteed.
+For image text input, considering that the image resolution is larger than the Channel in the early stage, we choose a smaller $SW$ in the early stage, and because the size of the input image text changes dynamically during processing, and the length and width resolution needs to be guaranteed.
 Divisible by $SW$, we empirically set $SW$ for stages using CSWin Self-Attention to $1$ or $2$ by default.
 Global Attention.
-As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual relationships between all character components.
+As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components.
 Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters.
 In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components.
-In addition, in order to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information.
+In addition, to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information.
 The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
-Positional Embedding.
-Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back.
-Typical positional Embedding mechanisms include absolute positional encoding (APE)Vaswani et al.
-(2017), relative positional encoding (RPE) Liu et al.
-(2021) and CPE.
-APE and RPE are designed for a specific input size and are not friendly to varying input resolutions.
-CPE takes the feature as input and can generate the positional encoding for arbitrary input resolutions.
-Then the generated positional encoding will be added onto the input feature.
-Locally-enhanced Positional Encoding (LePE)Dong et al.
-(2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
-As mentioned above, the two types of hybrid blocks aim to extract complementary different features.
-In VIPTR, blocks are processed locally first and then globally.
-At each stage, Local and Global Attentions are recurrently applied multiple times for comprehensive feature extraction.
-Effect of the two kinds of blocks arrangements will be shown in next ablation experiments.
+As mentioned above, the hybrid blocks in VIPTR are designed to extract distinct yet complementary features.
+In this way, blocks are initially processed locally before being addressed globally.
+At each stage, Local and Global Attention are recurrently applied multiple times for comprehensive feature extraction.
+Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
 3.5 Subsampling
-In order to save computational costs and extract significant features layer by layer, we added a down-sampling operationDu et al.
+To save computational costs and extract significant features layer by layer, we added a down-sampling operation Du et al.
 (2022) on the height dimension after the first and third attention mixing stages of VIPTR.
-Using the features output from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively.
+Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively.
 We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$.
@@
 For example, $N$ is set to $37$ for English when implemented.
-3.7 Architecture Variants
+3.7 Vision Permutation and Architecture Variants
 Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2.
-Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, number of mixing blocks.
-By varying them, VIPTR architectures with different number of parameters can be obtained.
-In summary, we constructed four typical architectures, namely VIPTRv1-T(Tiny), VIPTRv2-T, VIPTRv1-L(Large) and VIPTRv2-B(Base).
+Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, and number of mixing blocks.
+By varying them, VIPTR architectures with different numbers of parameters can be obtained.
+In summary, we constructed four typical architectures, namely VIPTRv1-T (Tiny), VIPTRv2-T, VIPTRv1-L (Large), and VIPTRv2-B (Base).
 Its detailed configuration is shown in Table 2.
@@
 Model $[C_{0},C_{1},C_{2},C_{3}]$ $[N_{1},N_{2},N_{3},N_{4}]$ Heads Attention Permutation Params (M) FLOPs (G)
-VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 3.99 0.26
+VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
 VIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
-VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.26 0.19
+VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
 VIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
-Table 2: VIPTR architecture variants (w/o counting the rectification module and linear classifier).
-4 Experiments
-4.1 Datasets
-For English STR task, our models are trained on two commonly used synthetic scene text datasets and a real scene dataset, i.e., MJSynth (MJ)Jaderberg et al.
+Table 2: VIPTR architecture variants (w/o counting the rectification module and CTC decoder).
+4 Experiments and Discussion
+Datasets
+Our English STR models are trained on two synthetic datasets, MJSynth (MJ) Jaderberg et al.
 (2014, 2016) and SynthText (ST) Gupta et al.
-(2016) trainset (72k+55k), COCO-Text Gomez et al.
-(2017)(38062) and the trainsets union of ICDAR 2013 (IC13)Karatzas et al.
-(2013), ICDAR 2015 (IC15)Karatzas et al.
-(2015), IIIT5K Mishra et al.
-(2012) and Street View Text (SVT) Wang et al.
+(2016) trainsets (72k+55k), along with real-world data from COCO-Text Gomez et al.
 (2017).
-Then the models are tested on six public benchmarks as follows: IC13 testset contains 857 selected regular testing images.
-SVT testset has 647 testing images cropped form Google Street View.
-IIIT5K-3000 is collected from the website and contains 3000 test images.
-IC15 is taken with Google Glasses without careful position and focusing, and testset we used has 1811 images.
-Street View Text Perspective (SVTP)Phan et al.
-(2013) is also cropped from Google Street View.
-There are 639 test images in this set.
-CUTE80 (CUTE) is proposed in Risnumawan et al.
-(2014) for curved text recognition.
-288 testing images are cropped from full images by using annotated words.
-For Chinese recognition task, we use the Chinese Scene and Document Dataset Yu et al.
+Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) Karatzas et al.
+(2013) testset with 857 regular images, Street View Text (SVT) Wang et al.
+(2011) testset with 647 images from Google Street View, IIIT5K-3000 Mishra et al.
+(2012) sourced from websites, ICDAR 2015 (IC15) Karatzas et al.
+(2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) Phan et al.
+(2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
+For the Chinese recognition task, we employ the Chinese Scene and Document Dataset Yu et al.
 (2021).
-The former is a public dataset containing 509,164, 63,645 and 63,646 training, validation, and test images.
-A total of 5,880 different characters appear.
-The latter is also a public dataset, which containing 400,000, 50,000 and 50,000 training, validation, and test images.
-A total of 4,865 different characters appear.
-The validation set is utilized to determine the best model, which is then assessed using the test set.
+The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters.
+The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters.
+Implementations
+The recommended training and testing configurations for VIPTR are listed in Table 3.
+When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$.
+Then, It was not used when training the Chinese model.
+All models are trained by PyTorch on four Tesla V100 GPUs.
+Setup English model Chinese model
+Rectification module Shi et al. (2018)
+Distortion correction $32\times 64$
+Optimizer AdamW
+Weight decay 0.05
+Lr scheduler Cosine annealing
+Input size $32\times 96$ $32\times 320$
+Batch size 1024 512
+Initial learning rate 5e-4×$\times\frac{batchsize}{2048}$ 3e-4×$\times\frac{batchsize}{512}$
+Training epochs 49 100
+Warm-up epochs 2 5
+Evaluation metric Word accuracy
+Table 3: Training configurations for VIPTR.
+Comparison with State-of-the-Art
+As indicated in Table 6, we compare VIPTR with prior studies on six English and one Chinese scene benchmarks.
+Even the fastest variant, VIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM).
+In English datasets, VIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15.
+Both VIPTRv1-L and VIPTRv2-B demonstrate overall accuracy comparable to recent studies Fang et al.
+(2021); Tang et al.
+(2021); Du et al.
+(2022); Lyu et al.
+(2022) while maintaining simplicity and faster execution.
+Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by Du et al.
+(2022).
+Encouragingly, VIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed.
+Other VIPTR variants also exhibit advanced Chinese recognition accuracy.
+These results are attributed to VIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
+4.1 Further Analysis
+Ablation Experiments
+To gain a deeper insight into VIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations.
+The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images over 500 epochs of training.
+In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths.
+The Chinese Document Dataset is used for training and validation.
+Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 1500 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 1500 pixels.
+For efficiency, all the experiments are carried out by using VIPTR-T without the rectification module and data augmentation.
 Positional Embedding IC13 IC15 FLOPs (G)
@@
 Table 5: Ablation studies on inputs of different lengths.
-4.2 Implementation Details
-VIPTR uses the rectification module Shi et al.
-(2018), where the image text is resized to $32\times 64$ for distortion correction.
-We use the AdamW optimizer with weight decay of $0.05$ for training.
-For English models, when CSWin is used by Local Attention, the input image size is set to $32\times 96$.
-Otherwise, it is set to $32\times 100$.
-The initial learning rate are set to $\frac{5}{10^{4}}\times\frac{batchsize}{2048}$.
-The cosine learning rate scheduler with $2$ epochs linear warm-up is used in all $49$ epochs.
-Data augmentation like stretch, perspective distortion, blur and Gaussian noise, are randomly performed during the training with probability $0.4$.
-The alphabet includes all case-insensitive alphanumerics.
-The maximum annotation length corresponding to the training sample is set to $25$.
-For Chinese models, the input size is set to $32\times 320$ and all images are resized or padded to this scale during training.
-The initial learning rate are set to $\frac{3}{10^{4}}\times\frac{batchsize}{512}$.
-The cosine learning rate scheduler with $5$ epochs linear warm-up is used in all $100$ epochs.
-Data augmentation was not used for training.
-Word accuracy is used as the evaluation metric.
-All models are trained by using 4 Tesla V100 GPUs on PyTorch.
-4.3 Ablation Study
-To better understand VIPTR, we perform controlled experiments on both IC13 (regular text) and IC15 (irregular text) under different settings.
-The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images and 500epochs of training.
-In addition, we also add aset of comparative experiments to verify the effectiveness of models on arbitrary lengths’ image texts.
-The Chinese Document Dataset is used for training and validation.
-Three test sets are prepared.
-Testset1 is the standard test set of Chinese Document Dataset, Testset2 (2367) is croped from real documents with image length in the range $[320,1500]$.
-Testset3 (638) gets from the same origin, whose image length is above 1500.
-For efficiency, all the experiments are carried out by using VIPTR-T without rectification module and data augmentation.
-Methods English regular English unregular Avg Chinese Params Speed $\downarrow$
-IC13 SVT IIIT5k IC15 SVTP CUTE Scene (M) (ms)
+Methods English Regular English Unregular Avg. Chinese Scene Params (M) Speed $\downarrow$ (ms)
+IC13 SVT IIIT5k IC15 SVTP CUTE
 CRNN Shi et al. (2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
@@
 ViTSTR Atienza (2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
-SVTR-TDu et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.03 4.5
+SVTR-T Du et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
 SVTR-L Du et al. (2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
 ASTER Shi et al. (2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
-NRTRSheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84 - 31.7 160
-SARLi et al. (2019) 91 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120
+NRTR Sheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
+SAR Li et al. (2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
 SATRN Lee et al. (2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
 SRN Yu et al. (2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
-ABINetFang et al. (2021) 97.4 93.5 96.2 86 89.3 89.2 92.6 - 36.7 51.3
-VisionLANWang et al. (2021b) 95.7 91.7 95.8 83.7 86 88.5 91.2 - 32.8 28
-VSTTang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64 -
+ABINet Fang et al. (2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
+VisionLAN Wang et al. (2021b) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
+VST Tang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64.0 -
 MATRN Na et al. (2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
-MaskOCR*Lyu et al. (2022) 97.7 93.7 95.4 86.6 89 87.5 92.5 68.6 36 24.1
-MaskOCRLyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100 -
-TrOCRLi et al. (2023b) 98.3 93.2 91 84 91 89.6 90.2 - 558 20.8
-AutoSTRZhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84 89.5 - 10.4 207
-PREN*Yan et al. (2021) 94.7 92 92.1 79.2 83.9 81.3 88 - 29.1 40
-VIPTRv1-T(Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.89 4.2
-VIPTRv2-T 95.8 92.6 92.6 85.4 86 85 90.3 65.8 5.1 3.3
+MaskOCR* Lyu et al. (2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
+MaskOCR Lyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
+TrOCR Li et al. (2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
+AutoSTR Zhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84.0 89.5 - 10.4 207.0
+PREN* Yan et al. (2021) 94.7 92.0 92.1 79.2 83.9 81.3 88.0 - 29.1 40.0
+VIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
+VIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
 VIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
-VIPTRv2-B(Base) 97 94.2 94.8 88 90 90.2 92.7 71 22.1 8.0
+VIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 71.0 22.1 8.0
 Table 6: Results on six English and Chinese scene benchmarks tested against existing methods.
 The model types from top to bottom are CTC-based, attention-based, LM-based, and others.
-* means lightweight version of the corresponding model.
-The speed is the average inference time on one NVIDIA Tesla V100 GPU with batchsize 4 (to align with Du et al.
+* means the lightweight version of the corresponding model.
+The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with Du et al.
 (2022)) for 3000 English image texts.
-4.3.1 The Effective Permutation of Attention Blocks
+Effective Permutation of Attention Blocks
 There are various ways to match the global and Local Attention blocks in each stage.
@@
 $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks.
-$[L_{i}][L_{j}//G_{s}][G_{t}]$ means in stage 2-3, Local and Global Attention blocks are carried in parallel.
+$[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel.
 The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA.
 Others are defined similarly.
-As can be seen, almost every scheme gives certain accuracy improvement compared to CRNN.
-We believe that the improvements are attributed to the comprehensive perception of character components features by self-attention mechanism.
-The relatively large gains on irregular text further explains the self-attention mixing block is helpful to feature learning in complex scenarios.
+It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN.
+We believe that the improvements are attributed to the comprehensive perception of character components features by the self-attention mechanism.
+The relatively large gains on irregular text further explain the self-attention mixing block is helpful to feature learning in complex scenarios.
 It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1.
@@
 It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
-4.3.2 The Effectiveness of Positional Embeddings.
-As seen in Table 3, different PEs behave slightly different in recognition accuracy.
-LePE scheme outperforms the default one (None) by 2.4% and 5.3% on the two datasets, indicating that it is effective especially for irregular text.
-4.3.3 The Results for different lengths inputs testing
-As seen in Table 4, the performance of the three models on the three testsets is quite different.
+4.1.1 Effectiveness of Positional Embeddings.
+As indicated in Table 4, various PE schemes exhibit slight variations in recognition accuracy.
+The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text.
+4.1.2 Results for different lengths inputs testing
+In Table 5, the performance of the three models on the three test sets is quite different.
 SVTR’s inference performance starts to collapse as the input length gets larger.
-VIPTR not only supports arbitrary length of image text input, but the inference performance will not decrease as the input length increases.
-4.4 Comparison with State-of-the-Art
-We compare VIPTR with previous works on six English benchmarks covering regular and irregular text and one Chinese scene dataset in Table 5.
-We first look into the result in English datasets.
-Even VIPTRv2-T, the smallest one of the VIPTR family, achieves highly competitive accuracy among the w/o LM ones.
-When further compared with LM-used ones and SVTR-L, VIPTRv1-L gains the best accuracy on IC15 among the six English benchmarks.
-VIPTRv1-L and VIPTRv2-B’s overall accuracy are on par with recent studiesFang et al.
-(2021); Tang et al.
-(2021); Du et al.
-(2022); Lyu et al.
-(2022).
-As a contrast, VIPTRv1-L and VIPTRv2-B enjoy its simplicity and runs faster.
-We then analyze the result on Chinese Scene Dataset.
-Du et al.
-(2022) gives the accuracy of 8 existing methods as shown in the table.
-Encouragingly, VIPTR performs considerably well.
-Compared with SVTR-L, the best one among the listed ones, accuracy gains 0.5%.
-Other VIPTR methods also have advanced Chinese recognition accuracy.
-The above results are explained as VIPTRs can better perceive multi-grained character component features.
-They also achieved breakthroughs in characterizing Chinese words with rich stroke patterns.
+VIPTR not only supports arbitrary lengths of image text input, but the inference performance will not decrease as the input length increases.
 5 Conclusion
-We have presented VIPTR, a customized vision permutable extractor for fast and efficient STR.
-It extracts multi-grained character features that describe both stroke-like local patterns and inter-component dependence of varying distance at multiple height scales.
-Therefore, enjoying the merits of accuracy, efficiency and cross-lingual versatility.
-VIPTR with different capacities are also devised to meet diverse application needs.
-Experiments on both English and Chinese benchmarks basically verify the proposed VIPTR.
-Highly competitive or even better accuracy is observed compared to state-of-the-art methods, while running faster.
-We hope to combine VIPTR with an efficient text decoder to further improve the performance of STR.
+In this study, we propose VIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR).
+VIPTR excels at extracting character features and capturing contextual dependencies between character components, all while being unrestricted by input size during inference.
+These grants VIPTR notable accuracy, efficiency, and cross-lingual adaptability.
+We have also developed different versions of VIPTR to cater to a range of application requirements.
+Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models.
+It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds.
+Our future goal is to integrate VIPTR with an efficient text decoder to further elevate STR performance.

### v5

@@
-VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition
-Abstract
-Scene Text Recognition (STR) is a challenging task that involves recognizing text within images of natural scenes.
-Although current state-of-the-art models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders.
-In this work, we propose the VIsion Permutable extractor for fast and efficient scene Text Recognition (VIPTR), which achieves an impressive balance between high-performance and rapid inference speeds in the domain of STR.
-Specifically, VIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by multiple self-attention layers, while eschewing the traditional sequence decoder.
-This design choice results in a lightweight and efficient model capable of handling inputs of varying sizes.
-Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of VIPTR.
-Notably, the VIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds.
-Meanwhile, the VIPTR-L (Large) variant attains greater recognition accuracy, while maintaining a low parameter count and favorable inference speed.
-Our proposed method provides a compelling solution for the STR challenge, which blends high accuracy with efficiency and greatly benefits real-world applications requiring fast and reliable text recognition.
+SVIPTR: Fast and Efficient Scene Text Recognition with Vision Permutable Extractor
+Abstract.
+Scene Text Recognition (STR) is an important and challenging upstream task for building structured information databases, which involves recognizing text within images of natural scenes.
+Although current state-of-the-art (SOTA) models for STR exhibit high performance, they typically suffer from low inference efficiency due to their reliance on hybrid architectures comprised of visual encoders and sequence decoders.
+In this work, we propose a VIsion Permutable extractor for fast and efficient Scene Text Recognition (SVIPTR), which achieves an impressive balance between high performance and rapid inference speeds in the domain of STR.
+Specifically, SVIPTR leverages a visual-semantic extractor with a pyramid structure, characterized by the Permutation and combination of local and global self-attention layers.
+This design results in a lightweight and efficient model and its inference is insensitive to input length.
+Extensive experimental results on various standard datasets for both Chinese and English scene text recognition validate the superiority of SVIPTR.
+Notably, the SVIPTR-T (Tiny) variant delivers highly competitive accuracy on par with other lightweight models and achieves SOTA inference speeds.
+Meanwhile, the SVIPTR-L (Large) attains SOTA accuracy in single-encoder-type models, while maintaining a low parameter count and favorable inference speed.
+Our proposed method provides a compelling solution for the STR challenge, which greatly benefits real-world applications requiring fast and efficient STR.
 The code is publicly available at https://github.com/cxfyxl/VIPTR.
-1 Introduction
-Text Recognition (TR) is crucial for multimodal document analysis, autonomous driving, office automation, and other applications Shi et al.
-(2016); Guo et al.
-(2022); Yang et al.
-(2020); Guo et al.
-(2023); Li et al.
-(2023a); Wu et al.
-(2023); Zhou et al.
-(2023).
+Scene Text Recognition, Attention mechanism, Vision Transformer, Visual-semantic analysis, Length-Insensitive
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679618isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Object recognitionccs: Computing methodologies Neural networksccs: Computing methodologies Scene understandingccs: Computing methodologies Image representations
+1. Introduction
+Text Recognition (TR) is crucial for multimodal document analysis and building structured databases, autonomous driving, office automation, and other applications (Shi et al., 2016; Guo et al., 2022; Yang et al., 2020; Guo et al., 2023; Li et al., 2023a; Wu et al., 2023; Zhou et al., 2023).
 Recent deep learning advancements have spurred growth in Scene Text Recognition (STR), which aims to digitize text from images in natural settings for use in language processing tasks.
@@
 Efforts are ongoing to improve recognition precision, with an emphasis on creating models that are accurate, fast, and lightweight for versatile deployment.
-Figure 1: Model Architecture Evolution for Scene Text Recognition.
-Scene text recognition initially relied on deep learning models like CRNNs Shi et al.
-(2016) that combined visual feature extraction with sequence processing, favored for their fast inference despite sensitivity to text disturbances.
-Subsequent models adopted encoder-decoder, autoregressive approaches, enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds.
-More recent advances involve visual semantic parsing frameworks that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise.
-Recent researchers 2021 (2020); Zhang et al.
-(2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, in addition, align the CE loss and CTC loss to improve the convergence speed of the CTC loss.
-From the perspective of simplifying the model architecture, some models used only CNN Borisyuk et al.
-(2018) or ViT Atienza (2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods.
-Inspired by the pyramid structure of Swin Transformer Liu et al.
-(2021), the SVTR model Du et al.
-(2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information.
-SVTR has four structural variants with different capacities.
+STR initially relied on deep learning models like CRNNs (Shi et al., 2016) that combined visual feature extraction with sequence processing (as shown in Figure 1(a)), favored for their fast inference despite sensitivity to text disturbances.
+Subsequent models adopted encoder-decoder, autoregressive approaches (as shown in Figure 1(b)), enhancing accuracy through contextual understanding but at the cost of slower word-by-word translation speeds.
+More recent advances involve visual semantic parsing frameworks (as shown in Figure 1(c)) that integrate semantic knowledge to improve recognition accuracy in specific contexts, which often require large models or intricate semantic systems, leading to efficiency compromise.
+Recent researchers (Hu et al., 2020; Zhang et al., 2023) tried to develop simple and effective model structures by using the attention mechanism to aggregate the sequence features corresponding to the same character, or regularising the extracted sequence frame by frame for supervised learning of individual characters.
+These strategies are used to align the Cross-Entropy (CE) loss and Connectionist Temporal Classification (CTC) loss, on the one hand, the convergence speed of CTC loss is improved without affecting inference speed, and on the other hand, the character and sequence features are strengthened.
+From the perspective of simplifying the model architecture, some models used only CNN (Borisyuk et al., 2018) or ViT (Atienza, 2021) for identification, which is efficient in inference but has a large gap in accuracy compared to SOTA methods.
+Recent researchers (Wang et al., 2022; Cheng et al., 2023; Fujitake, 2024) have proposed a series of innovative optimization methods for feature sequence decoders, simplifying the encoder-decoder architecture and effectively promoting a balance between high performance and inference efficiency.
+In addition, inspired by the pyramid structure of Swin Transformer (Liu et al., 2021), the SVTR model (Du et al., 2022), designed for single visual modality input, is proposed to simultaneously process the implicit local visual information in text images and unsupervised mining of global contextual semantic information.
 The STR models based on vision Transformer set a typical example for fast and effective STR schemes for a single visual modality.
 On one hand, it practices the effectiveness of replacing convolution with a self-attention mechanism to realize a new visual feature extractor.
-On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs.
-Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The actual inference time consumption of the model in general scenarios is still larger than that of CRNN with the same parameter magnitude.
-To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose VIPTR—a vision-based approach for a faster, more accurate, and cross-language universal STR model.
-Inspired by Cross-Shaped Windows Vision Transformer (CSwin) Dong et al.
-(2022) and other visual models based on enhanced self-attention mechanisms, VIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background.
+On the other hand, it reveals the importance of modeling global dependencies of lateral symbol sequences in text image recognition tasks, which even replace bidirectional LSTMs (Sun et al., 2022) and sequence decoders.
+Despite the outstanding performance of current ViT-based STR, it is clear that it still has several shortcomings: (1) The computation of self-attention relies on absolute position encoding and fixed-size mask operation, which limits the length of the input image; (2) In each stage, only one type of global and local relations is mined separately, and the distance between the steps of global dependencies is limited to save reasoning time, so the recognition effect of long text images is not good; (3) The Resource occupancy during training and CPU inference time consumption of models in general scenarios are still larger than that of CRNN with the same parameter magnitude.
+Figure 1. Model Architecture Evolution for Scene Text Recognition.
+[…]
+Model Architecture Evolution for Scene Text Recognition.
+To further enhance the vision model and meticulously control computational complexity to achieve more efficient STR while maintaining high accuracy, we propose SVIPTR—a visual-semantic-based approach for a faster, more accurate, and cross-language universal STR model.
+Inspired by Cross-Shaped Windows Vision Transformer (CSwin) (Dong et al., 2022) and other visual models based on enhanced self-attention mechanisms, SVIPTR first divides the text image evenly into multiple 2D patches, each patch contains part of the characters or background.
 This way of dividing patches facilitates the simultaneous extraction of local features of characters and the capture of potential contextual semantic dependencies between different characters.
-In each token mixing attention module, we try to use conditional positional embedding (CPE) Chu et al.
-(2021) combined with Manhattan self-attention (MaSA) Fan et al.
-(2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) Lou et al.
-(2023) module to model global dependencies.
+In each token mixing attention module, we try to use conditional positional embedding (CPE) (Chu et al., 2021) combined with Manhattan self-attention (MaSA) (Fan et al., 2023) and CSWin Attention to explore the local information of characters, and use Overlapping Spatial Reduction Attention (OSRA) (Lou et al., 2023) module to model global dependencies.
 Specifically, along the horizontal direction Model the global dependency between characters in the axial direction, model the global relationship from top to bottom of the same character along the vertical axis, and align to achieve weight sharing.
@@
 At the same time, we also use a pyramid structure for multi-stage local and global information hybrid modeling, which greatly improves the feature extraction capabilities of text images.
-To sum up, VIPTR can surpass the recognition effect and inference efficiency of SVTR at a smaller model size.
-The main contributions of this work can be summarized as.
-  • We verified that the single-vision model based on the self-attention mechanism can still achieve comparable accuracy to the high-level vision-language model in the scene text recognition task.
-At the same time, the use of sparse operators and different attention combinations can accelerate the calculation of the self-attention mechanism, and achieve the balance between performance and speed advantages that STR has been pursuing.
-  • We proposed VIPTR, a feature extraction module tailored for parsing text.
-The module can adapt to and accurately recognize the text image input of any length and has the versatility of cross-language recognition, which has a good application prospect.
-  • We verified the superiority of VIPTR on cross-language benchmark datasets and manually annotated industrial application datasets.
-Among them, VIPTRv1-L achieved SOTA in recognizing Chinese and English scene text.
-VIPTRv2-T achieves the most efficient inference while ensuring accuracy.
+To sum up, SVIPTR can ensure that the recognition effect and inference efficiency exceed the advanced vision Transformer-based recognition model at a smaller model scale.
+The main contributions of this work can be summarized as:
+  • We verify that the use of sparse operators and different Permutations and combinations of self-attention mechanisms can accelerate the calculation of Vision Transformer, and the model built based on these mechanisms can still achieve comparable accuracy to high-level vision-language models in STR tasks.
+Integrated, achieving the balance between performance and speed advantages that STR has been pursuing.
+  • We proposed SVIPTR, a visual-semantic feature extraction model tailored for parsing image text.
+The model can accurately recognize cross-language image text input and is insensitive to the length of image input, which has a good application prospect.
+  • We verified the superiority of SVIPTR on cross-language benchmark datasets and manually annotated industrial application datasets.
+Among them, SVIPTRv1-L achieved higher accuracy than other ViT-based encoder models on both Chinese and English STR.
+SVIPTRv1-L combined with BiLSTM decoder achieves the SOTA level on STR benchmarks.
+SVIPTRv2-T achieves the most efficient inference while ensuring accuracy.
 It is inferred on an NVIDIA V100 GPU with a parameter of 5.1M and consumes an average of 3.3ms per text image.
-2 Related Work
-Figure 2: Overall architecture of the proposed VIPTR.
+2. Related Work
+Figure 2.
+Overall architecture of the proposed SVIPTR.
 It is a four-stage network with three-stage height progressively decreasing.
 In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation.
-In particular, VIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel.
+In particular, SVIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel.
 At last, the recognition is conducted by the CTC decoder.
-CNN-Based Vision Modules
+[…]
+Overall architecture of the proposed SVIPTR.
+It is a four-stage network with three-stage height progressively decreasing.
+In each stage, a series of attention-mixing blocks are carried out and followed by a subsampling or combining operation.
+In particular, SVIPTR designs two visual feature fusion modes in vision permutation blocks: series and parallel.
+At last, the recognition is conducted by the CTC decoder.
+2.1. CNN-Based Vision Modules.
 Convolutional neural networks (CNN) serve as pivotal tools for extracting visual information in previous tasks.
-VGG Simonyan and Zisserman (2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model.
-Later, ResNet He et al.
-(2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance.
+VGG (Simonyan and Zisserman, 2014) demonstrated that a series of 3x3 convolutions is sufficient to train a high-performance model.
+Later, ResNet (He et al., 2016) introduced residual connections along with the batch normalization (BN) layer, which enabled the training of exceedingly deep neural networks and further improved performance.
 In the realm of STR, the CNN-based vision model transforms a pyramid structure capable of processing local features across multiple stages.
@@
 Therefore, these techniques continue to be utilized in our proposed STR architecture.
-Transformer-Based Vision Modules
-ViT Dosovitskiy et al.
-(2020) is the first work to introduce Transformer Vaswani et al.
-(2017); Yang et al.
-(2022); Touvron et al.
-(2021) into the construction backbone of vision tasks.
+2.2. Transformer-Based Vision Modules.
+ViT (Dosovitskiy et al., 2020) is the first work to introduce Transformer (Vaswani et al., 2017; Yang et al., 2022; Touvron et al., 2021) into the construction backbone of vision tasks.
 Experiments demonstrate that, with sufficient training data, ViT outperforms traditional CNNs in vision tasks.
-DeiT Touvron et al.
-(2021) introduces a new training strategy, enabling ViT to achieve superior performance on small-scale training sets such as ImageNet-1K.
+DeiT (Touvron et al., 2021) introduces a new training strategy, enabling ViT to achieve superior performance compared to CNNs on small-scale training sets such as ImageNet-1K.
 These methods demonstrate the efficacy of global dependency modeling through self-attention mechanisms.
@@
 Vision Transformers are integrated into STR tasks as visual feature encoders, leading to a significant improvement in accuracy.
-ViTSTR Atienza (2021), visionLAN Wang et al.
-(2021b), and MaskOCR Lyu et al.
-(2022) have all come close to the SOTA effect.
-SVTR-T both ensures high performance and creates the fastest recognition efficiency.
-Although the self-attention mechanism has greatly improved the effect of text recognition, its computational complexity is quadratic compared to the number of input tokens.
-As the resolution of the input image increases, the inference time consumption also increases quadratically.
-Efficient Self-Attention Based Vision Modules
-A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive Vaswani et al.
-(2017); Yang et al.
-(2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining.
+VisionLAN (Wang et al., 2021a), SVTR, MGP-STR (Wang et al., 2022) and MaskOCR (Lyu et al., 2022) have all come close to the SOTA effect.
+2.3. Sparse Self-attention and Spatial Modeling.
+A challenging problem in the design of Transformer is that the computational cost of global self-attention is very expensive (Vaswani et al., 2017; Yang et al., 2021, 2023), while local self-attention usually limits the interaction field of each token, resulting in insufficient information mining.
 The CSWin Self-Attention mechanism uses parallel computing to form the self-attention of the horizontal and vertical stripes of the cross-shaped window, achieving powerful feature modeling capabilities while limiting computing costs.
 Decomposed MaSA achieves efficient and significant feature representation by defining a spatial attention attenuation matrix and its decomposition paradigm based on two-dimensional Manhattan distance, effectively layering local information on a global scale.
-Spatial reduction attention (SRA) Wang et al.
-(2021a) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) Lou et al.
-(2023) into SRA to achieve better performance by using larger and overlapping patches.
-3 Method
-3.1 Design Guidelines
+Spatial reduction attention (SRA) (Wang et al., 2021b) effectively extracts global information by utilizing sparse token-region dependencies, and OSRA introduces overlapping spatial reduction (OSR) (Lou et al., 2023) into SRA to achieve better performance by using larger and overlapping patches.
+3. Method
+3.1. Design Guidelines
 In this study, we use some important properties of CNNs, ResNet, and several self-attention mechanisms in a new architecture inspired by hierarchical vision Transformers to design a paradigm of high-performance networks for STR.
 Here are the design guidelines we followed:
-  1. 1.
-    Explicitly inject character features into the front of the network in the form of locality bias.
-  2. 2.
-    Design a new permutation method to effectively fuse the character component features of the Local and Global Attention modules in the middle layer.
-  3. 3.
-    Implicitly handle image aspect and width resolution in the network initialization structure, canceling or replacing all modules that depend on the input size.
-3.2 Overall Architecture
-The schematic representation of the proposed VIPTR is depicted in Figure 2.
+  1. (1)
+    Explicitly vectorizing the character features in the form of segmented chunks and injecting them into the front of the network.
+  2. (2)
+    Design permutation and combination modes of self-attention mechanism to effectively fuse visual and semantic information of characters embeddings in the middle layer.
+  3. (3)
+    Implicitly handling the length information of input images during network initialization and designing shape-insensitive self-attention mechanisms to replace all modules that depend on the input size.
+3.2. Overall Architecture
+The schematic representation of the proposed SVIPTR is depicted in Figure 2.
 It is a four-stage network with three height progressively decreased layers.
@@
 Since each patch is associated with a text character or part of the background in the image, these patches are called character components.
-Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by downsampling or global average pooling operations along the height dimension.
-Both Local and Global Attention blocks are employed to concentrate on character feature extraction and capture dependencies between character components.
-The extractor serves to characterize character component features at different distances and multi-scales and the contextual dependencies between character components, and a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$ is generated, which captures multi-granularity character features.
-Finally, the CTC decoder is used to perform linear prediction and deduplication on the character sequence to obtain the recognition result.
-3.3 Patch Embedding
+Subsequently, four stages of feature extraction are performed at different scales, each stage consists of a series of self-attention-based hybrid blocks, followed by dimension reduction in half using convolutional or global average pooling operations along the height dimension.
+Finally, the extractor generated a representation called $B$ with a size of $1\times\frac{W}{4}\times C_{3}$, which captures multi-granularity characters’ sequence features.
+At the same time, the CTC decoder is used to linearly predict $B$ to obtain the recognition result.
+3.3. Patch Embedding
 The first step of the vision Transformer is to cut the image into several equal-sized patches, and then hand them over to the subsequent self-attention module for processing in the form of several sets of sequences.
@@
 Alternatively, we implement the patch embedding by using two consecutive $3\times 3$ convolutions with stride $2$ and BNs.
-This scheme Du et al.
-(2022) gradually increases feature dimensions, which is beneficial to feature fusion.
-Figure 3: Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
-3.4 Character and String Modeling
+This scheme (Du et al., 2022) gradually increases feature dimensions, which is beneficial to feature fusion.
+3.4. Character and String Modeling
 Existing studies mainly use feature sequences with height $1$ to represent image text.
@@
 However, these segments are usually noisy, especially for irregular text, failing to describe the characters of an image well.
-In contrast, the Vision Transformer introduces a two-dimensional feature representation, which enables more accurate discrimination between different characters and backgrounds.
-This suggests that two-dimensional features can be used to effectively represent the local strokes of characters and the potential contextual dependencies of strings.
-More specifically, image text requires features to be extracted from two perspectives.
-The first involves local character features, including elements like strokes and fonts, representing internal associations of a character and morphological features.
-The second is context relations, such as correlations between different characters or between text and non-text components.
-Therefore, we design two classes of attention-mixing blocks with different receptive fields to sense relevance.
-Mathematically speaking, for a sequence of character components $B_{i-1}B_{i}$ in the previous stage, it is first reshaped as a sequence of features.
-Upon the input of the attention mixing block, Layer Normalization (LN) and Position Embedding (PE) are first applied.
+In contrast, vision Transformer introduces two-dimensional feature representation, which not only enables more accurate discrimination between different characters and backgrounds but also fully exploits the potential contextual dependencies in patch sequence.
+Figure 3.
+Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
+[…]
+Illustration of (a) Cross-Shaped Windows Self-Attention (CSWin), (b) Decomposed Manhattan Self-Attention (D-MaSA), (c) Multi-Head Self-Attention (MHSA), and (d) Overlapping Spatial Reduction Attention (OSRA).
+More specifically, image text requires complementary features to be extracted from two perspectives.
+The first involves local visual features, representing internal associations and morphological features of each character.
+The second is contextual semantic relations, such as correlations between different characters or between text and non-text components.
+Therefore, we design two classes of self-attention blocks with different receptive fields to sense relevance.
+Mathematically speaking, a sequence of character components $B_{i-1}B_{i}$ in the previous stage, is first reshaped as a sequence of features.
+Upon inputting the self-attention block, Layer Normalization (LN) and Position Embedding (PE) are first applied.
 Subsequently, the self-attention mechanism is employed for dependency modeling.
@@
 Since self-attention is permutation invariant and ignores tokens’ positional information, positional Embedding (PE) is widely used in Transformers to add such information back.
-Typical positional Embedding mechanisms include absolute positional encoding (APE) Vaswani et al.
-(2017), relative positional encoding (RPE) Liu et al.
-(2021), and CPE.
+Typical positional Embedding mechanisms include absolute positional encoding (APE) (Vaswani et al., 2017), relative positional encoding (RPE) (Liu et al., 2021), and CPE.
 APE and RPE are designed for a specific input size and are not friendly to varying input resolutions.
@@
 Then the generated positional encoding will be added to the input feature.
-Locally-enhanced Positional Encoding (LePE) Dong et al.
-(2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
+Locally-enhanced Positional Encoding (LePE) (Dong et al., 2022) shares a similar spirit as CPE, but proposes to add the positional encoding as a parallel module to the self-attention operation and operates on projected values in each Transformer block.
 Local Attention.
@@
 Drawing on the design ideas of MaSA and CSWin Self-Attention, we designed two efficient self-attention mechanisms that explicitly model local features on a two-dimensional coordinate system.
-MaSA is shown in Figure 3LABEL:sub@subfig:dmsa.
+MaSA is shown in Figure 3dLABEL:sub@subfig:dmsa.
 The spatial attenuation matrix of the image means that character features are highlighted globally and background features are weakened.
-CSWin Self-Attention is shown in Figure 3LABEL:sub@subfig:cswin.
+CSWin Self-Attention is shown in Figure 3dLABEL:sub@subfig:cswin.
 This mechanism is implemented by executing local self-attention of horizontal and vertical stripes of step size called Split-Window ($SW$) decomposed from a certain cross window in parallel.
@@
 Global Attention.
-As shown in Figure 3LABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components.
+As shown in Figure 3dLABEL:sub@subfig:mhsa, Global Attention calculates the contextual dependencies between all character components.
 Since text and non-text are the two main elements in an image, this universal blend can create long-distance dependencies between the components of different characters.
 In addition, using the self-attention mechanism can also weaken the influence of non-text components while enhancing the importance of text components.
-In addition, to save computational costs, we introduced OSRA, as Figure 3LABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information.
+In addition, to save computational costs, we introduced OSRA, as Figure 3dLABEL:sub@subfig:osra, which adds overlapping convolution operations to MHSA to sparse character components in some channels into more salient features without losing edge information.
 The use of salient features greatly reduces the number of tokens required to calculate the self-attention mechanism, thus improving computational efficiency and performance.
-As mentioned above, the hybrid blocks in VIPTR are designed to extract distinct yet complementary features.
-In this way, blocks are initially processed locally before being addressed globally.
-At each stage, Local and Global Attention are recurrently applied multiple times for comprehensive feature extraction.
-Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
-3.5 Subsampling
-To save computational costs and extract significant features layer by layer, we added a down-sampling operation Du et al.
-(2022) on the height dimension after the first and third attention mixing stages of VIPTR.
+3.5. Height Dimension Reduction
+To save computational costs and extract significant features layer by layer, we added the Height Dimension Reduction operation (HDR) after each self-attention block in every stage of SVIPTR, like classic CNN backbones (Borisyuk et al., 2018).
 Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively.
-We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$.
-The Subsampling block halves the height while maintaining a constant width.
-It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension.
-We also increase the channel dimension $C_{i}$ to compensate for the information loss.
-3.6 Combining and Prediction
-In the final stage, we use combinatorial operations Du et al.
-(2022) to integrate sequence features.
-It first uses AdaptiveAvgPool to compress the height dimension to $1$, followed by fully connected layers, Hardswish activation, and dropout.
+In the previous stage, we then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$.
+In the final stage, we first use AdaptiveAvgPool to compress the height dimension to $1$, followed by Fully Connected Layers, Hardswish Activation, and Dropout.
 By doing so, the character components are further compressed into feature sequences, where each element is represented by a feature of length $C_{3}$.
-Using AdaptiveAvgPool avoids applying convolution to very small embeddings of one dimension, such as height $2$, while being able to apply a variety of initial input heights.
+Using AdaptiveAvgPool avoids applying convolution to embeddings with very small dimensions, such as height $2$, while being able to apply a variety of initial input heights.
+The HDR block halves the height while maintaining a constant width.
+It not only reduces the computational cost but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension.
+We also increase the channel dimension $C_{i}$ to compensate for the loss of information.
+3.6. Prediction
 We finally use the CTC decoder to perform parallel linear prediction on the obtained combined features to achieve recognition.
@@
 For example, $N$ is set to $37$ for English when implemented.
-3.7 Vision Permutation and Architecture Variants
-Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2.
-Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, and number of mixing blocks.
-By varying them, VIPTR architectures with different numbers of parameters can be obtained.
-In summary, we constructed four typical architectures, namely VIPTRv1-T (Tiny), VIPTRv2-T, VIPTRv1-L (Large), and VIPTRv2-B (Base).
+3.7. Vision Permutation and Architecture Variants
+There are various ways to permutate the global and Local Attention blocks in each stage.
+In previous researchs (Du et al., 2022), it was found that placing the Local Attention block before the Global Attention block is beneficial to both feature mining of individual symbols and to guiding the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model.
+As shown in Figure 2, in the 2-3 stages of SVIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called SVIPTRv1 and SVIPTRv2.
+Symbolically, $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks.
+$[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel.
+The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA.
+Others are defined similarly.
+Effect of the two kinds of blocks Permutations can be seen in ablation experiments.
+We design several hyperparameters in SVIPTR, including channel depth and number of heads per stage, and number of mixing blocks.
+By varying them, SVIPTR architectures with different numbers of parameters can be obtained.
+In summary, we constructed four typical architectures, namely SVIPTRv1-T (Tiny), SVIPTRv2-T, SVIPTRv1-L (Large), and SVIPTRv2-B (Base).
 Its detailed configuration is shown in Table 3.
+Table 1. Ablation study on Attention Blocks’ permutations.
 Atten. Permutation IC13 IC15 FLOPs (G)
-CRNN (None) 75.3 56.7 0.69
-SVTR-T 77.9 62.5 0.29
-$[L_{1}][L_{1}G_{1}][G_{1}]$ 79.1 65 0.269
-$[L_{1}][L_{1}G_{2}][G_{1}]$ 79.1 65.2 0.258
-$[L_{2}][L_{2}G_{1}][G_{1}]$ 79 65.1 0.282
-$[L_{2}][L_{2}G_{2}][G_{1}]$ 79.3 65.3 0.27
-$[L_{1}][L_{2}G_{1}][G_{1}]$ 78.8 64.9 0.27
-$[L_{2}][L_{1}G_{1}][G_{1}]$ 78.9 65.1 0.27
-$[L_{1}][L_{2}G_{2}][G_{1}]$ 78.7 65 0.258
-$[L_{2}][L_{1}G_{2}][G_{1}]$ 78.8 64.8 0.258
-$[L_{1}][L_{1}//G_{1}][G_{1}]$ 75.5 60.5 0.195
-$[L_{1}][L_{1}//G_{2}][G_{1}]$ 75.4 60.5 0.192
-$[L_{2}][L_{2}//G_{1}][G_{1}]$ 75.6 60.4 0.197
-$[L_{2}][L_{2}//G_{2}][G_{1}]$ 75.8 61 0.194
-$[L_{1}][L_{2}//G_{1}][G_{1}]$ 75 59.9 0.196
-$[L_{2}][L_{1}//G_{1}][G_{1}]$ 75.2 59.7 0.196
-$[L_{1}][L_{2}//G_{2}][G_{1}]$ 74.9 60.1 0.193
-$[L_{2}][L_{1}//G_{2}][G_{1}]$ 75.2 60.3 0.193
-Table 1: Ablation study on Attention Blocks’ permutations.
-Model $[C_{0},C_{1},C_{2},C_{3}]$ $[N_{1},N_{2},N_{3},N_{4}]$ Heads Attention Permutation Params (M) FLOPs (G)
-VIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
-VIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
-VIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
-VIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
-Table 2: VIPTR architecture variants (w/o counting the rectification module and CTC decoder).
-4 Experiments and Discussion
+CRNN (None) 91.1 69.4 0.69
+SVTR-T 93.5 74.8 0.29
+$[L_{1}][L_{1}G_{1}][G_{1}]$ 94.4 80 0.269
+$[L_{1}][L_{1}G_{2}][G_{1}]$ 94.4 80.2 0.258
+$[L_{2}][L_{2}G_{1}][G_{1}]$ 94.3 80.1 0.282
+$[L_{2}][L_{2}G_{2}][G_{1}]$ 94.6 80.4 0.27
+$[L_{1}][L_{2}G_{1}][G_{1}]$ 94.1 79.9 0.27
+$[L_{2}][L_{1}G_{1}][G_{1}]$ 94.2 80.1 0.27
+$[L_{1}][L_{2}G_{2}][G_{1}]$ 94 80 0.258
+$[L_{2}][L_{1}G_{2}][G_{1}]$ 94.1 79.8 0.258
+$[L_{1}][L_{1}//G_{1}][G_{1}]$ 92.8 75.6 0.195
+$[L_{1}][L_{1}//G_{2}][G_{1}]$ 92.7 75.6 0.192
+$[L_{2}][L_{2}//G_{1}][G_{1}]$ 92.9 75.5 0.197
+$[L_{2}][L_{2}//G_{2}][G_{1}]$ 93.1 76.1 0.194
+$[L_{1}][L_{2}//G_{1}][G_{1}]$ 92.3 75 0.196
+$[L_{2}][L_{1}//G_{1}][G_{1}]$ 92.5 74.8 0.196
+$[L_{1}][L_{2}//G_{2}][G_{1}]$ 92.2 75.2 0.193
+$[L_{2}][L_{1}//G_{2}][G_{1}]$ 92.5 75.4 0.193
+4. Experiments and Discussion
+4.1. Experiment Settings
 Datasets
-Our English STR models are trained on two synthetic datasets, MJSynth (MJ) Jaderberg et al.
-(2014, 2016) and SynthText (ST) Gupta et al.
-(2016) trainsets (72k+55k), along with real-world data from COCO-Text Gomez et al.
-(2017).
-Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) Karatzas et al.
-(2013) testset with 857 regular images, Street View Text (SVT) Wang et al.
-(2011) testset with 647 images from Google Street View, IIIT5K-3000 Mishra et al.
-(2012) sourced from websites, ICDAR 2015 (IC15) Karatzas et al.
-(2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) Phan et al.
-(2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
-For the Chinese recognition task, we employ the Chinese Scene and Document Dataset Yu et al.
-(2021).
+Our English STR models are trained on two synthetic datasets, MJSynth (MJ) (Jaderberg et al., 2014, 2016) and SynthText (ST) (Gupta et al., 2016) trainsets (72k+55k), along with real-world data from COCO-Text (Gomez et al., 2017).
+Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) (Karatzas et al., 2013) testset with 857 regular images, Street View Text (SVT) (Wang et al., 2011) testset with 647 images from Google Street View, IIIT5K-3000 (Mishra et al., 2012) sourced from websites, ICDAR 2015 (IC15) (Karatzas et al., 2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) (Phan et al., 2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
+For the Chinese recognition task, we employ the Chinese Scene and Document Dataset (Yu et al., 2021).
 The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters.
 The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters.
-Implementations
-The recommended training and testing configurations for VIPTR are listed in Table 3.
+We use the validation set to select the best model, which is subsequently evaluated on the test set.
+Implementations.
+The recommended training and testing configurations for SVIPTR are listed in Table 2.
 When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$.
@@
 All models are trained by PyTorch on four Tesla V100 GPUs.
+Table 2. Training configurations for SVIPTR.
 Setup English model Chinese model
-Rectification module Shi et al. (2018)
+Rectification module (Shi et al., 2018)
 Distortion correction $32\times 64$
@@
 Evaluation metric Word accuracy
-Table 3: Training configurations for VIPTR.
-Comparison with State-of-the-Art
-As indicated in Table 6, we compare VIPTR with prior studies on six English and one Chinese scene benchmarks.
-Even the fastest variant, VIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM).
-In English datasets, VIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15.
-Both VIPTRv1-L and VIPTRv2-B demonstrate overall accuracy comparable to recent studies Fang et al.
-(2021); Tang et al.
-(2021); Du et al.
-(2022); Lyu et al.
-(2022) while maintaining simplicity and faster execution.
-Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by Du et al.
-(2022).
-Encouragingly, VIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed.
-Other VIPTR variants also exhibit advanced Chinese recognition accuracy.
-These results are attributed to VIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
-4.1 Further Analysis
-Ablation Experiments
-To gain a deeper insight into VIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations.
-The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images over 500 epochs of training.
+Comparison with State-of-the-Art.
+As indicated in Table 4, we compare SVIPTR with prior studies on six English and one Chinese scene benchmarks.
+Even the fastest variant, SVIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM).
+In English datasets, SVIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15.
+Both SVIPTRv1-L and SVIPTRv2-B demonstrate overall accuracy comparable to recent studies (Fang et al., 2021; Tang et al., 2021; Lyu et al., 2022) while maintaining simplicity and faster execution.
+Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by (Du et al., 2022).
+Encouragingly, SVIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed.
+Other SVIPTR variants also exhibit advanced Chinese recognition accuracy.
+These results are attributed to SVIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
+Table 3. SVIPTR architecture variants (w/o counting the rectification module and CTC decoder).
+Model $[C_{0},C_{1},C_{2},C_{3}]$ $[N_{1},N_{2},N_{3},N_{4}]$ Heads Attention Permutation Params (M) FLOPs (G)
+SVIPTRv1-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}G_{2}][G_{1}]$ 4.0 0.26
+SVIPTRv1-L $[192,256,512,384]$ $[3,7,2,9]$ $[6,8,8,16]$ $[L_{2}][L_{2}G_{2}][G_{1}]$ 37.7 2.31
+SVIPTRv2-T $[64,128,256,192]$ $[3,3,3,3]$ $[2,4,4,8]$ $[L_{1}][L_{1}//G_{2}][G_{1}]$ 3.2 0.19
+SVIPTRv2-B $[128,256,384,256]$ $[3,6,6,9]$ $[4,8,8,12]$ $[L_{2}][L_{2}//G_{2}][G_{1}]$ 20.2 1.18
+Table 4.
+Results on six English and Chinese scene benchmarks tested against existing methods.
+The model types from top to bottom are CTC-based, attention-based, LM-based, and using dedicated decoders.
+* means the lightweight version of the corresponding model.
+The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with (Du et al., 2022)) for 3000 English image texts.
+Methods English Regular English Unregular Avg. Chinese Scene Params (M) Speed $\downarrow$ (ms)
+IC13 SVT IIIT5k IC15 SVTP CUTE
+CRNN (Shi et al., 2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
+Rosetta (Borisyuk et al., 2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
+ViTSTR (Atienza, 2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
+SVTR-T (Du et al., 2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
+SVTR-L (Du et al., 2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
+ASTER (Shi et al., 2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
+NRTR (Sheng et al., 2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
+SAR (Li et al., 2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
+SATRN (Lee et al., 2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
+SRN (Yu et al., 2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
+ABINet (Fang et al., 2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
+VisionLAN (Wang et al., 2021a) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
+MATRN (Na et al., 2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
+MaskOCR* (Lyu et al., 2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
+MaskOCR (Lyu et al., 2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
+TrOCR (Li et al., 2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
+MGP-STR-T (Wang et al., 2022) 96.5 93.2 96.3 86.2 89.4 90.6 92.7 - 5.4 10.6
+MGP-STR(Fuse) (Wang et al., 2022) 97.3 94.7 96.4 87.2 91.0 90.2 93.3 - 21.0 12.0
+SVTR-L+DCTC (Zhang et al., 2023) 97.4 93.7 96.9 87.3 88.5 92.3 93.3 73.9 40.8 18.0
+SVIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
+SVIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
+SVIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
+SVIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 72.0 22.1 8.0
+SVIPTRv1-L+BiLSTM+DCTC 98.0 94.8 96.1 88.8 90.4 91.2 93.6 74.2 45.8 22.0
+4.2. Ablation Study
+To gain a deeper insight into SVIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations.
+The training set and experimental Settings in the English scenario are unchanged.
 In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths.
 The Chinese Document Dataset is used for training and validation.
-Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 1500 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 1500 pixels.
-For efficiency, all the experiments are carried out by using VIPTR-T without the rectification module and data augmentation.
-Positional Embedding IC13 IC15 FLOPs (G)
-None 76.7 59.7 0.25
-APE 77.3 63.5 0.26
-CPE 77.3 63.6 0.26
-LePE 79.1 65 0.26
-Table 4: Ablation study on Positional Embeddings.
-Model Testset1 Testset2 Testset3
-CRNN 97.5 91 91.6
-SVTR-T 98.2 47.5 5.9
-VIPTRv1-T 98.8 92.1 93.7
-Table 5: Ablation studies on inputs of different lengths.
-Methods English Regular English Unregular Avg. Chinese Scene Params (M) Speed $\downarrow$ (ms)
-IC13 SVT IIIT5k IC15 SVTP CUTE
-CRNN Shi et al. (2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
-Rosetta Borisyuk et al. (2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
-ViTSTR Atienza (2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
-SVTR-T Du et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
-SVTR-L Du et al. (2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
-ASTER Shi et al. (2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
-NRTR Sheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
-SAR Li et al. (2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
-SATRN Lee et al. (2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
-SRN Yu et al. (2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
-ABINet Fang et al. (2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
-VisionLAN Wang et al. (2021b) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
-VST Tang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64.0 -
-MATRN Na et al. (2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
-MaskOCR* Lyu et al. (2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
-MaskOCR Lyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
-TrOCR Li et al. (2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
-AutoSTR Zhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84.0 89.5 - 10.4 207.0
-PREN* Yan et al. (2021) 94.7 92.0 92.1 79.2 83.9 81.3 88.0 - 29.1 40.0
-VIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
-VIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
-VIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
-VIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 71.0 22.1 8.0
-Table 6: Results on six English and Chinese scene benchmarks tested against existing methods.
-The model types from top to bottom are CTC-based, attention-based, LM-based, and others.
-* means the lightweight version of the corresponding model.
-The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with Du et al.
-(2022)) for 3000 English image texts.
-Effective Permutation of Attention Blocks
-There are various ways to match the global and Local Attention blocks in each stage.
-In previous researchv Du et al.
-(2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model.
-The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations.
-$[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks.
-$[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel.
-The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA.
-Others are defined similarly.
+Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 640 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 640 pixels.
+For efficiency, all the experiments are carried out by using SVIPTR-T without the rectification module and data augmentation.
+Effective Permutation of Attention Blocks.
+The top half and the bottom half of Table 1 respectively show the performance of SVIPTRv1-T and SVIPTRv2-T composed of various Local and Global Attention and their permutations.
 It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN.
@@
 It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
-4.1.1 Effectiveness of Positional Embeddings.
+In general, SVIPTR has better performance than SVTR without using rectification module, and it is the fastest and highest performance visual feature encoder.
+Effectiveness of Positional Embeddings.
 As indicated in Table 5, various PE schemes exhibit slight variations in recognition accuracy.
 The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text.
-4.1.2 Results for different lengths inputs testing
+The use of LePE not only improves the recognition accuracy without affecting the inference efficiency but also is insensitive to the input length, which enhances the versatility of SVIPTR.
+Table 5. Ablation study on Positional Embeddings.
+Model + Positional Embedding IC13 IC15 FLOPs (G)
+SVIPTRv1-T + None 93.7 79.7 0.25
+SVIPTRv1-T + APE 94 78.2 0.26
+SVIPTRv1-T + CPE 94.3 78.9 0.26
+SVIPTRv1-T + LePE 94.6 80.3 0.26
+Results for different lengths inputs testing.
 In Table 6, the performance of the three models on the three test sets is quite different.
 SVTR’s inference performance starts to collapse as the input length gets larger.
-VIPTR not only supports arbitrary lengths of image text input, but the inference performance will not decrease as the input length increases.
-5 Conclusion
-In this study, we propose VIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR).
-VIPTR excels at extracting character features and capturing contextual dependencies between character components, all while being unrestricted by input size during inference.
-These grants VIPTR notable accuracy, efficiency, and cross-lingual adaptability.
-We have also developed different versions of VIPTR to cater to a range of application requirements.
+SVTR uses APE and SWin self-attention with static Windows, which does not have CNN-like translation invariance.
+So the length of inference images are limited by the size of training images.
+If inference images’ lengths are beyond the pre-set, resizing the image is required, which results in distortion.
+SVIPTR is not only length-insensitive, but the inference performance will not decrease as the input length increases.
+Table 6. Ablation studies on inputs of different lengths.
+Model Testset1 Testset2 Testset3
+CRNN 97.5 91 91.6
+SVTR-T 98.2 47.5 5.9
+SVIPTRv1-T 98.8 92.1 93.7
+4.3. Attention Visualization
+We have visualized the attention maps of SVIPTR when decoding different character components.
+Finally, the proposed SVIPTRv2-T model without TPS and STN (It is obtained by training a large number of image texts with a uniform shape size of 32x96) with the weakest performance is selected to generate the attention map and shown in this paper, which further proves the high performance of all variants of SVIPTR.
+As shown in Figure 4, the picture in the first row is the original image (the shape size is $32\times 480$), and the right side is the model prediction result.
+The following attention map can be interpreted as the global and local attention mechanisms that play a role in the whole recognition process.
+The first attention map represents the degree of attention given to the character ”e” in the recognition process.
+From here, it can be seen that the model can maintain strong attention to the target character in a relatively long time span, which verifies the effectiveness of the global attention mechanism.
+The subsequent attention map is composed of selecting an appropriate timestamp corresponding to the attention of each character component in the original image.
+From the phenomenon that almost each character is covered by a bright color, it can be seen that the model allocates enough attention to the local vicinity of each character component, which verifies the effectiveness of the local attention mechanism.
+Moreover, the attention does not significantly decay as the character sequence becomes longer, which means that the dependencies between different character components are successfully captured, which further indicates that the model has good long-distance reasoning ability.
+Figure 4 indicates that SVIPTRv2-T captures multiple features and cues locally, throughout, and across characters, which is consistent with the claim that SVIPTR senses multi-granularity character component features.
+It also verifies that the inference of SVIPTR is not sensitive to the length of the image, and the weakest performing variant also has strong long-distance inference ability.
+This again illustrates the effectiveness of SVIPTR and its variants.
+Figure 4. Visualization of the attention maps for SVIPTRv2-T. The character in red means it is missed or an error.
+[…]
+Visualization of the attention maps for SVIPTRv2-T. The character in red means it is missed or an error.
+5. Conclusion
+In this study, we propose SVIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR).
+SVIPTR excels at extracting and fusing visual and semantic features of image text, and it is insensitive to the length of the inference input.
+These grants SVIPTR notable accuracy, efficiency, and cross-lingual adaptability.
+We have also developed different versions of SVIPTR to cater to a range of application requirements.
 Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models.
 It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds.
-Our future goal is to integrate VIPTR with an efficient text decoder to further elevate STR performance.
+Our future goal is to integrate SVIPTR with an efficient text decoder to further elevate STR performance.
+Acknowledgements.
+This work was supported in part by the National Natural Science Foundation of China (Grant Nos.
+U1636211, U2333205, 61672081, 62302025, 62276017), a fund project: State Grid Co., Ltd.
+Technology R&D Project (ProjectName: Research on Key Technologies of Data Scenario-based Security Governance and Emergency Blocking in Power Monitoring System, Proiect No.: 5108-202303439A-3-2-ZN), the 2022 CCF-NSFOCUS Kun-Peng Scientific Research Fund and the Opening Project of Shanghai Trusted Industrial Control Platform and the State Key Laboratory of Complex & Critical Software Environment (Grant No.
+SKLSDE-2021ZX-18).